OpenStudy (anonymous):

how about m-4 - 3m-1 =1 --- ---- 3 5

jhonyy9 (jhonyy9):

m-4 3m-1 ---- - ----- = 1 5(m-4)-3(3m-1)=15 3 5 5m-20-9m+3 =15 -4m=32 m=32/-4 m=-8

OpenStudy (anonymous):

[5(m-4)]-[3(3m-1)]/15{LCM} = 1 [(5m-20)-(9m-3)]/15 = 1 5m-20-9m+3 = 1*15 -4m-17 = 15 -4m = 15+17 m = 32/-4 m = -8